OpenStudy (anonymous):

What is the equation, in slope-intercept form, of the line that passes through (0, 5) and has a slope of −1?

OpenStudy (anonymous):

Slope intercept form. y = mx + b. So, it has a slope of m = -1. Plug it in! y = -1x + b Replace x, y with the point you were given. 5 = -1(0) + b 5 = b The final equation is y = -1x + 5